OpenStudy (anonymous):

Identify the equation or coordinates of the following conic and its equation. Identify the points where the ellipse (x^2/25)+(y^2/13)=1 intersects the x-axis.

OpenStudy (lgbasallote):

no more square roots huh?

OpenStudy (anonymous):

\[\frac{x^2}{25} + \frac{y^2}{13} =1\] & Nope, thankfully, those are over ! New section..